# CAD-Editor: A Locate-then-Infill Framework with Automated Training Data Synthesis for Text-Based CAD Editing

### v1

@@ 6.3 Ablation Studies
Method JSD $\downarrow$ D-CLIP $\uparrow$ VR $\uparrow$
CAD-Editor-mini w/ Basic 0.78 - 0.12 88.8
CAD-Editor-mini w/ Step 0.70 - 0.07 90.1
CAD-Editor w/ Direct 0.77 - 0.16 86.1
CAD-Editor w/ L-I 0.67 0.21 96.5
CAD-Editor w/ L-I & HS 0.65 0.27 95.6
@@
Selective Dataset. In Section 5.2, we propose improving the performance of the infilling stage with selective data curated with human annotation. We conduct experiments both with or without such human selection, denoted as CAD-Editor w/ L-I & HS and CAD-Editor w/ L-I. Table 2 shows that this strategy results in the greatest improvement in D-CLIP, highlighting its effectiveness in aligning editing instructions with the edited CAD model.
@@ 8 Conclusion
Appendix A Additional Results
We present a qualitative comparison between the directly fine-tuned LLM and our Locate-then-Infill framework in Figure 10. Compared to the directly fine-tuned approach, our framework enhances generation quality, improves text-CAD alignment and ensures greater output stability.
Further qualitative comparisons between CAD-Editor and baseline methods are presented in Figure 11, highlighting the superior performance of CAD-Editor across diverse editing conditions.
@@ Appendix B Stepwise Captioning Strategy
For stepwise captioning, we employ GPT-4o four times per CAD pair to generate the final editing instruction at the image level, while LLaMA-3-70B is used for the sequence level. The detailed prompt is shown in Figure 12.
Figure 12: The detailed prompt for stepwise captioning.
Moreover, we provide a comparison between basic captioning and our stepwise captioning results in Figure 13.
Figure 13: Comparisons between basic captioning method and our stepwise captioning method.
@@ Appendix C Locate-then-Infill Framework
We use task-specific prompts for locating task and infilling task, with details shown in Table 3.
@@
Table 3: Prompt for Locate-then-Infill framework.
Appendix D Baseline Prompting
Here, we provide the detailed prompt for the baseline method in Figure 14.
Figure 14: Prompt for GPT-4o-IC.

### v2

@@ 6.4 Ablation Studies
Selective Dataset. In Section 5.2, we propose improving the performance of the infilling stage with selective data curated with human annotation. We conduct experiments both with or without such human selection, denoted as CAD-Editor w/ L-I & HS and CAD-Editor w/ L-I. Table 2 shows that this strategy results in the greatest improvement in both JSD and D-CLIP scores, demonstrating its effectiveness in enhancing generation quality and better aligning editing instructions with the edited CAD model.
@@ 8 Conclusion
Acknowledgements
We thank Ruiyu Wang for his assistance in conducting one of the baseline experiments. We also appreciate the constructive comments from the anonymous reviewers.
@@ Acknowledgements
Appendix A CAD-Editor Dataset Preprocessing
The DeepCAD dataset exhibits a severe data imbalance. Among deduplicated 137,012 training models, over 91.1% have three or fewer SE operations, while only 6.2% have 4–5 SEs. Models with more than 5 SEs make up less than 3%, and those exceeding 10 SEs account for just 0.4%. To mitigate the impact of this long-tail distribution, we limit our dataset to models with at most 3 SE pairs.
Additionally, the design variation generation model can sometimes introduce noise by altering a CAD model too drastically or not at all. To ensure data quality, we implement a data filtering strategy. We filter out CAD pairs with too many changes by excluding examples with more than three editing instructions or more than five <mask> tokens. We also filter out pairs with no significant changes by excluding instructions containing phrases like “no transformation is needed”.
@@ Appendix B Stepwise Captioning Strategy
To implement stepwise captioning, we utilize GPT-4o four times per CAD pair to generate the final image-level editing instructions. For sequence-level, we employ LLaMA-3-70B. The detailed prompts are illustrated in Figure 11.
Figure 11: Detailed prompt used for stepwise captioning.
Furthermore, Figure 12 presents a comparison between the basic captioning approach and our proposed stepwise captioning method.
Figure 12: Comparison between the basic captioning method and our stepwise captioning method.
@@ Appendix C Locate-then-Infill Framework
We adopt task-specific prompts for both the locating and infilling stages. The detailed prompt designs for each task are shown in Table 3.
Table 3: Prompt for Locate-then-Infill framework.
@@
Appendix D Additional Qualitative Results
We present qualitative comparisons between the directly fine-tuned LLM and our proposed Locate-then-Infill framework in Figure 13. Compared to the direct fine-tuning approach, our framework improves generation quality, enhances text-CAD alignment, and ensures greater output stability.
Additional qualitative comparisons between CAD-Editor and other baseline methods are shown in Figure 14, demonstrating the superior performance of CAD-Editor under various editing conditions.
Appendix E Scaling & Generalization
Due to the severe data imbalance in DeepCAD, we constrain the SE length during training. To evaluate the scalability of our method, we conduct an additional experiment in which we artificially balance the dataset across different SE lengths. Specifically, we construct a training set containing 4,000 samples for each SE length from 1 to 5, ensuring a uniform distribution. The model is then evaluated on separate test sets corresponding to each SE length. As shown in Table 4, CAD-Editor consistently outperforms the baselines across all SE lengths. Model performance declines slightly as SE length increases. Since generating longer sequences is inherently more challenging, this result demonstrates that our method generalizes well to complex CAD structures given sufficient training data.
To further assess generalization, we evaluate CAD-Editor’s generalization by testing a model trained on DeepCAD directly on Fusion 360 dataset (Willis et al., 2021). As shown in Table 5 , CAD-Editor outperforms baselines, confirming its generalization ability to datasets with different shape distributions.
We present additional results generated by CAD-Editor in Figure 15, including cases with more than three SEs and examples sourced from the Fusion 360 dataset.
Table 4: Quantitative evaluation across different numbers of sketch-extrude (SE) operations. We construct a dataset of 20,000 examples with a uniform distribution over SE counts from 1 to 5. JSD and D-CLIP values are scaled by $10^{2}$. $\uparrow$: higher is better, $\downarrow$: lower is better.
Method JSD $\downarrow$ D-CLIP $\uparrow$ VR $\uparrow$
1 2 3 4 5 1 2 3 4 5 1 2 3 4 5
GPT-4o-Basic 5.98 2.86 3.68 3.82 3.95 -0.31 -0.90 -0.61 -0.47 -0.97 56.5 64.2 59.4 61.5 58.3
GPT-4o-IC 5.42 1.75 2.29 2.40 3.67 0.39 -0.06 -0.50 -0.11 -0.38 76.2 85.0 76.5 75.5 73.1
CAD-Editor 4.08 1.46 2.01 2.35 3.15 0.41 -0.23 -0.06 -0.02 -0.29 84.3 91.6 87.5 79.3 79.5
Table 5: Quantitative evaluation on the dataset generated from the Fusion 360 Gallery. Lower JSD and higher D-CLIP and VR values indicate better performance. The results demonstrate that our method exhibits strong generalization ability.
[…]
GPT4o-Basic 4.58 -0.45 57.7
GPT4o-IC 2.59 -0.39 81.5
CAD-Editor 2.34 0.41 93.9
Figure 15: Extended results from CAD-Editor on more complex scenarios, including examples with more than three SEs and examples sourced from Fusion 360 datasets.
Appendix F Baseline Editing Methods
This section presents the detailed prompt used in the GPT-4o-IC baseline, as illustrated in Figure 16. The GPT-4o-Basic prompt follows the same format, excluding the in-context examples.
Figure 16: Prompt employed in baseline methods for CAD model editing.